# spacing=nonfrench

### v2

@@ 2 Preliminaries
Let $L_{r}$ be the leader of round $r$. Each replica keeps the highest voted round $r_{vote}$, the highest locked round $r_{lock}$  [^2] Corresponds to the preferred round in [36]., current round number $r_{cur}$, the highest quorum certificate $qc_{high}$, and the last-round timeout certificate $TC_{last}$. The current view number $v_{cur}$ is not used and remains 00 throughout. Replicas initialize $r_{vote}=0$, $r_{lock}=0$, $r_{cur}=1$, $qc_{high}$ as the QC of the genesis block of round 00, $TC_{last}=\bot$ and enter round $1$. Steady State Protocol for Replica $i$ Propose. Upon entering round $r$, the leader $L_{r}$ multicasts a block $B=[id,qc_{high},r,v_{cur},txn]$. Vote. Upon receiving the first proposal $B=[id,qc,r,v,txn]$ from $L_{r}$ in round $r$, execute Advance Round, Lock, and then Commit (defined below). If $r=r_{cur}$, $v=v_{cur}$, $r>r_{vote}$ and $qc.r\geq r_{lock}$, vote for $B$ by sending the threshold signature share $\{id,r,v\}_{i}$ to $L_{r+1}$, and update $r_{vote}\leftarrow r$. Lock. ($2$-chain lock rule) Upon observing any $qc$ [^3] May be formed from votes (by a leader) or contained in a proposal or a timeout messsage., let $qc^{\prime}$ be the QC contained in the block certified by $qc$ (i.e., $qc^{\prime}$ is the parent of $qc$), the replica updates $r_{lock}\leftarrow\max(r_{lock},qc^{\prime}.r)$, and $qc_{high}\leftarrow\max(qc_{high},qc)$. Commit. ($3$-chain commit rule) Whenever there exist three adjacent certified blocks $B,B^{\prime},B^{\prime\prime}$ in the chain with consecutive round numbers, i.e., $B^{\prime\prime}.r=B^{\prime}.r+1=B.r+2$, the replica commits $B$ and all its ancestors. Pacemaker Protocol for Replica $i$ Advance Round. If the replica receives or forms a quorum certificate $qc$ of round $r\geq r_{cur}$, it updates its current round $r_{cur}\leftarrow r+1$ and $TC_{last}\leftarrow\bot$. If the replica receives or forms a timeout certificate $tc$ of round $r\geq r_{cur}$, it updates its current round $r_{cur}\leftarrow r+1$ and $TC_{last}\leftarrow tc$. Timer and Timeout. Upon entering round $r$, the replica sends the round-$(r-1)$ TC to $L_{r}$ if it has the TC, and resets its timer to count down for a predefined time interval (timeout $\tau$). When the timer expires, the replica stops voting for round $r_{cur}$ and multicasts a timeout message $\langle\{r_{cur}\}_{i},qc_{high},TC_{last}\rangle_{i}$ where $\{r_{cur}\}_{i}$ is a threshold signature share. Upon receiving a timeout message or TC, execute Advance Round, Lock, and then Commit. Upon receiving $f+1$ timeouts of a same round $r\geq r_{cur}$, the replica stops voting for round $r_{cur}$ and multicasts a timeout message $\langle\{r_{cur}\}_{i},qc_{high},TC_{last}\rangle_{i}$ where $\{r_{cur}\}_{i}$ is a threshold signature share. Upon receiving $12f+$ 1 timeouts, form a TC.
@@ 2.1 Terminology
  • Timeout Certificate. A timeout message of round $r$ by a replica contains the replica’s threshold signature share on $r$, its $qc_{high}$, and a $TC$ of round $r-1$ if $qc_{high}.round\neq r-1$ (otherise $TC=\bot$). A timeout certificate (TC) is formed by a quorum of $n-f=2f+1$ timeout messages, containing a threshold signature on a round number $r$ produced by combining $12f+$ 1 signature shares $\{r\}$ from the timeout messages, and $12f+$ 1 $qc_{high}$’s. A valid TC should only contain high-QCs with round numbers $<TC.r$, and this will be checked implicitly when a replica receives a TC.
@@ 3 Jolteon Design
Replicas keep the same variables as DiemBFT in Figure 1. Steady State Protocol for Replica $i$ Changes from DiemBFT in Figure 1 are marked in blue. Propose. Upon entering round $r$, the leader $L_{r}$ multicasts a block $B=[id,qc_{high},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}tc},r,v_{cur},txn]$, where $tc=tc_{r-1}$ if $L_{r}$ enters round $r$ by receiving a round-$(r-1)$ $tc_{r-1}$, and $tc=\bot$ otherwise. Vote. Upon receiving the first proposal $B=[id,qc,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}tc},r,v,txn]$ from $L_{r}$, execute Advance Round, Lock, and then Commit (defined below). If $r=r_{cur}$, $v=v_{cur}$, $r>r_{vote}$ and ( (1) $r=qc.r+1$, or (2) $r=tc.r+1$ and $qc.r\geq\max\{qc_{high}.r~{}|~{}qc_{high}\in tc\}$), vote for $B$ by sending the threshold signature share $\{id,r,v\}_{i}$ to $L_{r+1}$, and update $r_{vote}\leftarrow r$. Lock. ($1$-chain lock rule) Upon seeing a $qc$ (formed by votes or contained in proposal or timeouts), the replica updates $qc_{high}\leftarrow\max(qc_{high},qc)$. Commit. ($2$-chain commit rule) Whenever there exists two adjacent certified blocks $B,B^{\prime}$ in the chain with consecutive round numbers, i.e., $B^{\prime}.r=B.r+1$, the replica commits $B$ and all its ancestors. Pacemaker Protocol for Replica $i$ Identical to DiemBFT in Figure 1.
@@ Acknowledgments
We thank Nibesh Shrestha, Isaac Doidge and Raghavendra Ramesh from Supra for pointing out an issue in the description of DiemBFT Pacemarker in the previous version of the paper. The description is fixed in this version.

### v3

@@ 2 Preliminaries
Let $L_{r}$ be the leader of round $r$. Each replica keeps the highest voted round $r_{vote}$, the highest locked round $r_{lock}$  [^2] Corresponds to the preferred round in [36]., current round number $r_{cur}$, the highest quorum certificate $qc_{high}$, and the last-round timeout certificate $TC_{last}$. The current view number $v_{cur}$ is not used and remains 00 throughout. Replicas initialize $r_{vote}=0$, $r_{lock}=0$, $r_{cur}=1$, $qc_{high}$ as the QC of the genesis block of round 00, $TC_{last}=\bot$ and enter round $1$. Steady State Protocol for Replica $i$ Propose. Upon entering round $r$, the leader $L_{r}$ multicasts a block $B=[id,qc_{high},r,v_{cur},txn]$. Vote. Upon receiving the first proposal $B=[id,qc,r,v,txn]$ from $L_{r}$ in round $r$, execute Advance Round, Lock, and then Commit (defined below). If $r=r_{cur}$, $v=v_{cur}$, $r>r_{vote}$ and $qc.r\geq r_{lock}$, vote for $B$ by sending the threshold signature share $\{id,r,v\}_{i}$ to $L_{r+1}$, and update $r_{vote}\leftarrow r$. Lock. ($2$-chain lock rule) Upon observing any $qc$ [^3] May be formed from votes (by a leader) or contained in a proposal or a timeout messsage., let $qc^{\prime}$ be the QC contained in the block certified by $qc$ (i.e., $qc^{\prime}$ is the parent of $qc$), the replica updates $r_{lock}\leftarrow\max(r_{lock},qc^{\prime}.r)$, and $qc_{high}\leftarrow\max(qc_{high},qc)$. Commit. ($3$-chain commit rule) Whenever there exist three adjacent certified blocks $B,B^{\prime},B^{\prime\prime}$ in the chain with consecutive round numbers, i.e., $B^{\prime\prime}.r=B^{\prime}.r+1=B.r+2$, the replica commits $B$ and all its ancestors. Pacemaker Protocol for Replica $i$ Advance Round. If the replica receives or forms a quorum certificate $qc$ of round $r\geq r_{cur}$, it updates its current round $r_{cur}\leftarrow r+1$ and $TC_{last}\leftarrow\bot$. If the replica receives or forms a timeout certificate $tc$ of round $r\geq r_{cur}$, it updates its current round $r_{cur}\leftarrow r+1$ and $TC_{last}\leftarrow tc$. Timer and Timeout. Upon entering round $r$, the replica sends the round-$(r-1)$ TC to $L_{r}$ if it has the TC, and resets its timer to count down for a predefined time interval (timeout $\tau$). When the timer expires, the replica stops voting for round $r_{cur}$ and multicasts a timeout message $[\langle r_{cur},qc_{high}\rangle_{i},TC_{last}]$. Upon receiving a timeout message or TC, execute Advance Round, Lock, and then Commit. Upon receiving $f+1$ timeouts of a same round $r\geq r_{cur}$, the replica stops voting for round $r_{cur}$ and multicasts a timeout message $[\langle r_{cur},qc_{high}\rangle_{i},TC_{last}]$. Upon receiving $12f+$ 1 timeouts, form a TC.
@@ 2.1 Terminology
  • Timeout Certificate. A timeout message of round $r$ by a replica contains a signed $(r,qc_{high})$ message, and a $TC$ of round $r-1$ if $qc_{high}.round\neq r-1$ (otherwise $TC=\bot$), i.e., $[\langle r,qc_{high}\rangle_{i},TC]$. A timeout certificate (TC) is formed by a quorum of $n-f=2f+1$ timeout messages, containing $12f+$ 1 signed (round, highest QC) messages from these timeout messages. A valid TC of round $r$ should only contain signed $(r,*)$, and only contain highest QCs with round numbers $<r$, and this will be checked implicitly when a replica receives a TC.
@@ 3 Jolteon Design
Replicas keep the same variables as DiemBFT in Figure 1. Steady State Protocol for Replica $i$ Changes from DiemBFT in Figure 1 are marked in blue. Propose. Upon entering round $r$, the leader $L_{r}$ multicasts a block $B=[id,qc_{high},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}tc},r,v_{cur},txn]$, where $tc=tc_{r-1}$ if $L_{r}$ enters round $r$ by receiving a round-$(r-1)$ $tc_{r-1}$, and $tc=\bot$ otherwise. Vote. Upon receiving the first proposal $B=[id,qc,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}tc},r,v,txn]$ from $L_{r}$, execute Advance Round, Lock, and then Commit (defined below). If $r=r_{cur}$, $v=v_{cur}$, $r>r_{vote}$ and ( (1) $r=qc.r+1$, or (2) $r=tc.r+1$ and $qc.r\geq\max\{qc_{high}.r~{}|~{}qc_{high}\in tc\}$), vote for $B$ by sending the threshold signature share $\{id,r,v\}_{i}$ to $L_{r+1}$, and update $r_{vote}\leftarrow r$. Lock. ($1$-chain lock rule) Upon seeing a $qc$ (formed by votes or contained in proposal or timeouts), the replica updates $qc_{high}\leftarrow\max(qc_{high},qc)$. Commit. ($2$-chain commit rule) Whenever there exists two adjacent certified blocks $B,B^{\prime}$ in the chain with consecutive round numbers, i.e., $B^{\prime}.r=B.r+1$, the replica commits $B$ and all its ancestors. Pacemaker Protocol for Replica $i$ Identical to DiemBFT in Figure 1.
@@ Acknowledgments
We thank Nibesh Shrestha, Isaac Doidge and Raghavendra Ramesh from Supra for pointing out an issue in the description of DiemBFT Pacemarker in the previous version of the paper. We thank Matthieu Rambaud for pointing out an issue in the DiemBFT timeout certificate definition in the previous version of the paper. The issues are fixed in this version.